zaydenjones:

christian randomly selects students from his grade to rate a math test as easy moderate or difficult. of all the students he surveyed 13 said the test was easy 11 rated it as moderate and 3 found it difficult. assuming that all the students took the test how many of the 162 students would rate the test as something other than easy?

AZ:

So he took a survey and he got the results: 13 rated it as easy 11 rated it as moderate 3 rated it as difficult so he's surveyed a total of 13 + 11 + 3 = 27 students

AZ:

We're trying to find out how many of the 162 students would rate the test as something OTHER THAN easy so that means the students had to rate it as either moderate or difficult so first, of our 27 students surveyed, 13 said it was easy. That means the rest rated it as something other than easy how many students would that be? basically what is 27-13 =

zaydenjones:

14

AZ:

Good! Yes!

AZ:

So now we know that 14 students out of the 27 said that the exam was something other than easy now we set up a proportion to see how many would say it out of the 162 students

AZ:

We would get  we have to solve for x to get our answer

AZ:

Do you know how to solve for x?

zaydenjones:

Would you do the butterfly method?

AZ:

I have not heard it being called that but it sounds correct. We would cross multiply

1 attachment
zaydenjones:

Yes. That is what we call it.

AZ:

Yes! Then do that and 'x' will be how many students out of 162 would rate it as something other than easy

zaydenjones:

so, 14 x 162 / 27 = 84

AZ:

And that's your answer! Well done!

zaydenjones:

Awesome!!! Thank you for explaining it so well. It was so easy to do.

AZ:

Of course! I'm glad I could help :)